OpenStudy (anonymous):

tom,ted,tony,and terry worked together on a job.the job paid according to the amount of work he contributed to the job.ted and tony earned $280 together.ted and tom have $260,and ted and terry got $220.what percent of the job did each boy do

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@mathgeek27

OpenStudy (anonymous):

@mathgeek27

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@mathgeek27

OpenStudy (anonymous):

i know is 760, but I think a piece of info is missing

OpenStudy (anonymous):

*all money added together is 760

OpenStudy (anonymous):

my teacher gave me an equation but i dont know how to find the percetage of each boy do you want me to give you the equation

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

he said t= ted (The constant) t+(280-t)+(260-t)+(220-t)=500 t-t-t-t+280+220+260=500 the two ts cancel out so its 2t+760=500 280 is ted and tony 260 is ted and tom and 220 is ted and terry just in case you wanted to know the other t the first t is ted and the other t in the parenthesis is his friends

OpenStudy (mathmath333):

first of all ur question is incomplete

OpenStudy (anonymous):

personally, I'm stumped, but i'll get my friend to help u, He's great:)

OpenStudy (anonymous):

ok mathgeek27

OpenStudy (mathmath333):

Tom, Ted, Tony and Terry worked together on a job. The job paid $500. Each boy was paid according to the amount of work he contributed to the job. Ted and Tony earned $280 together. Ted and Tom have $260 and Ted and Terry got $220. What percent of the job did each boy do this is the correct one

OpenStudy (anonymous):

@-StudyShark-

OpenStudy (anonymous):

sorry i couldn't be of further help :(

OpenStudy (anonymous):

Good luck :)

OpenStudy (anonymous):

its fine @mathgeek27

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

do you think you can help

OpenStudy (mathmath333):

wait

OpenStudy (anonymous):

@-StudyShark-

OpenStudy (anonymous):

@phi

OpenStudy (mathmath333):

\(\large\tt \color{black}{tom+ted+tony+ terry=500}\) \(\large\tt \color{black}{ted+tony=280}\) \(\large\tt \color{black}{tom+ted=260}\) \(\large\tt \color{black}{ted+ terry=220}\)

OpenStudy (anonymous):

but what percentage does each boy make

OpenStudy (phi):

this is a "system of equations" with 4 unknowns and 4 equations. all those "t names" are very confusing, so let me use w,x,y,z w= tom, x= ted, y= tony, z= terry w + x + y + z = 500 x +y = 280 w+ x = 260 x + z = 220

OpenStudy (anonymous):

so @phi how im i supposed find out what percentage each boy makes

OpenStudy (phi):

It still looks complicated. but we could use the last equation to solve for z: z = 220 - x and the second equation to solve for y: y= 280 -x and the 3rd equation to solve for w: w= 260 -x and use those in the first equation

OpenStudy (phi):

4 equations and 4 unknowns are hard. but to continue z = 220 - x y= 280 -x w= 260 -x w + x + y + z = 500 (260-x) + x + (280-x) + (220-x) = 500 260 + (-x +x) + 280 -x + 220 - x = 500 the -x+x =0 the numbers 260+280+220 add up to 760 the -x - x add up to -2x so we get 760 - 2x = 500 can you solve for x?

OpenStudy (anonymous):

yes i think i can solve for x @phi

OpenStudy (phi):

what do you get ?

OpenStudy (anonymous):

x equals 130 @phi

OpenStudy (phi):

w= tom, x= ted, y= tony, z= terry ok, x (ted) makes 130 total is 500 for percent use part/whole * 100 130/500 * 100

OpenStudy (anonymous):

so for example terry percentage will be 220/100

OpenStudy (anonymous):

i mean 220/500 *100 @phi

OpenStudy (phi):

what did you get for ted?

OpenStudy (anonymous):

130

OpenStudy (phi):

what percent ?

OpenStudy (anonymous):

26%

OpenStudy (anonymous):

because you said 130 over 500 multiply by 100

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes. now to find the others, we have to be careful. If you followed what we did then you know we have these equations: x= 130 z = 220 - x y= 280 -x w= 260 -x w= tom, x= ted, y= tony, z= terry

OpenStudy (anonymous):

so it will be 220-130=90 90/500 * 100 will be the percentage for z

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes, and z is terry

OpenStudy (anonymous):

oh and for terry i will do the same for each and ever one of them @phi

OpenStudy (anonymous):

but ted percent was 26 right

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes so far ted is 26 and terry (z) is 90/500 * 100 btw, I would multiply top and bottom by 2 to get 180/1000 * 100 and simplify that to 180/10 = 18%

OpenStudy (anonymous):

oh ok do you want me to do that with all of them or do it the way i was doing it @phi

OpenStudy (anonymous):

and thank you so much btw

OpenStudy (phi):

do it the long way. the other was a side note.

OpenStudy (anonymous):

oh thank you so much @phi

OpenStudy (phi):

as a check, when you are done, all 4 percentages should add up to 100%

OpenStudy (anonymous):

ok thank you @ phi